OpenStudy (anonymous):

What is the slope of the graph of 4x – 16y = 15? A.-4 B.-1/4 C.1/4 D.4 @hartnn

OpenStudy (domebotnos):

4x - 16y = 15 -4x -4x -16y = -4x + 15 _____ ____ ___ -16 -16 -16 y = 1/4x - 15/16 The answer is C.

OpenStudy (anonymous):

after this one i have 6 more

OpenStudy (anonymous):

Which ordered pair is a solution to the equation 3x + 5y = 9? A. (–2, –3) B. (–2, 3) C. (2, –3) D. (2, 3) @domebotnos

OpenStudy (domebotnos):

3(-2) + 5(3) = 9 -6 + 15 = 9 9 = 9 The answer is B.